OpenStudy (anonymous):

1. The Doe family is driving from Philadelphia to Boston. So far,they have driven 167 miles. This is 129 miles less than the total distance they must travel. How many miles is Philadelphia from Boston. 2. Jon has $1230 in his savings account. This is $400 more than he needs to buy a new TV. Write and solve an equation to find out how much the TV costs. 3.Jane earns $5.50 per hour babysitting. She wants to buy a CD player that costs $71.50, including tax.How many hours will she need to work to earn the money for the CD player? 4. A cat can jump the height of up to 5 times the length of

OpenStudy (anonymous):

...its tail. How high can it jump if its tail is 13 inches long?

OpenStudy (anonymous):

1. 167 = x - 129 x = 167 + 129 2. 1230 = x + 400 x = 1230 - 400

OpenStudy (anonymous):

3. 5.50x = 71.50 x = 71.50 / 5.50 4. x = 13 * 5

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

\[56-r=8\]